# Gapfree graphs and powers of edge ideals with linear quotients

### v1

@@ Abstract.
Let $G$ be a gapfree graph and let $I(G)$ be its edge ideal. An open conjecture of Nevo and Peeva states that $I(G)^{q}$ has a linear resolution for $q\gg 0$. We investigate a stronger conjecture that if $I(G)^{q}$ has linear quotients for some integer $q$, then $I(G)^{q+1}$ also has linear quotients. We give a partial solution to this conjecture. It is known that if $G$ does not contain a cricket, a diamond, or a $C_{4}$, then $I(G)^{q}$ has a linear resolution for $q\geq 2$. We construct a family of gapfree graphs $G$ containing cricket, diamond, $C_{4}$ and $C_{5}$ as induced subgraphs of $G$ for which $I(G)^{q}$ has linear quotients for $q\geq 2$.
@@
edge ideal, gapfree graph, linear quotients.
@@ 1. Introduction
Let $G$ be a finite simple graph, i.e., a graph that does not contain loops or multiple edges, with vertex set $V(G)=\{1,\ldots,n\}$ and edge set $E(G)$. Let $S=K[x_{1},\ldots,x_{n}]$ denote the polynomial ring in $n$ variables over a field $K$. The edge ideal of $G$, denoted by $I(G)$, is the ideal of $S$ generated by squarefree quadratic monomials $x_{i}x_{j}$ for all $\{i,j\}\in E(G)$ (cf. [12]). The study of edge ideals has become a central topic in combinatorial commutative algebra, showcasing the deep interplay between combinatorial methods and algebraic structures. The present paper builds upon this rich foundation, advancing the understanding of powers of edge ideals and their algebraic properties.
@@
Conjecture 1.1 has garnered significant interest (cf. [2, 5, 6, 10] and references therein), but remains unresolved in general. While it has been verified for $q=2,3$ when $I(G)$ has small regularity (see [10]) and for random graphs with restricted edge probability (see [3]), progress has been limited. A promising approach to this challenging conjecture is to explore the stronger property of having linear quotients. Specifically, we make the following conjecture:
@@
If $G$ is gapfree and $I(G)^{q}$ has linear quotients for some $q\in{\mathbb{N}}$, then $I(G)^{N}$ has linear quotients for all $N\geq q$.
@@
For instance, the pentagon $C_{5}$ is gapfree, but its complementary graph is not chordal. Consequently, $I(C_{5})$ does not have a linear resolution. However, $C_{5}$ does not contain any induced subgraph that is isomorphic to a cricket (or a diamond, or $C_{4}$), so $I(C_{5})^{q}$ have linear resolution for all $q\geq 2$. On the other hand, it was shown (cf. [8, Corollary 10.1.7 and Theorems 10.1.9 and 10.2.5]) that if the complementary graph of $G$ is chordal then all powers of $I(G)$ have linear quotients. Beside the anticycles, as exhibited in a recent preprint [4] which appears after the first draft of this paper was written, there are currently no other known examples of a gapfree graph $G$ whose complementary graph is not chordal, yet for which all powers $I(G)^{q}$ with $q\geq 2$ have linear quotients. This motivates the search for natural classes of such graphs.
@@
In this paper, we establish a framework for addressing Conjecture 1.2. In Section 2, we examine the duplication technique which uses duplication of vertices in $G$ to produce new graphs $G^{*}$. We show that if a power $I(G)^{q}$ has linear quotients, then so does $I(G^{*})^{q}$ for certain derived graphs $G^{*}$ (Proposition 2.2). In Section 5, a similar result is shown when the vertex duplication is replaced by a clique, as long as the resulting graph is still gapfree (Theorem 5.3).
@@
Section 3 studies Conjecture 3.1, which makes a strong case for Conjecture 1.2. Particularly, Proposition 3.2 shows that Conjecture 3.1 holds for the pentagon $C_{5}$. In other words, all powers $I(C_{5})^{q}$ with $q\geq 2$ have linear quotients. Section 4 extends the study in Section 3 to a special class of graphs which are gapfree but may contain critical induced subgraphs such as crickets, diamonds and $C_{4}$. In Propositions 4.2 and 4.3, it is shown that, for these graphs, all powers $I(G)^{q}$ with $q\geq 2$ have linear quotients. The main result of the section (Theorem 4.6) extends the aforementioned result to an infinite family of graphs.
@@
Finally, Section 6 presents a refined analysis of modified version of Conjectures 3.1 and 1.2, culminating in Theorem 6.4, which exhibits that if $I(G)^{2},\ldots,I(G)^{7}$ have linear quotients with a special ordering of their generators, then $I(G)^{q}$, for all $q\geq 2$, have linear quotients. This result reduces the verification of Conjectures 3.1 and 1.2 to finitely many powers and, therefore, suggests that resolving these conjectures may be within reach.
2. Duplication of finite graphs
Let $G$ be a finite graph on $V(G)=\{1,\ldots,n\}$ with $E(G)=\{e_{1},\ldots,e_{s}\}$ and $S=K[x_{1},\ldots,x_{n}]$ the polynomial ring in $n$ variables over a field $K$. For an edge $e=\{i,j\}\in E(G)$, we define $u_{e}=x_{i}x_{j}\in S$. Recall that the edge ideal of $G$ is the ideal $I(G)$ of $S$ which is generated by those monomials $u_{e}$ with $e\in E(G)$. To simplify the notation, unless misleading, we may write $e$ instead of $u_{e}$. For example, $e_{1}^{3},e_{1}e_{2}^{2}\in I(G)^{3}$.
@@ 2. Duplication of finite graphs
Let, in general, $I$ be an ideal of $S$ which is generated by monomials of the same degree and $\{u_{1},\ldots,u_{r}\}$ the unique minimal system of monomial generators of $I$. We then call $u_{1},\ldots,u_{r}$ the generators of $I$. We say that $I$ has linear quotients if there is an ordering $u_{1}>\cdots>u_{r}$ of the generators of $I$ for which the colon ideal $(u_{1},\ldots,u_{i}):u_{i+1}$ is generated by a subset of $\{x_{1},\ldots,x_{n}\}$ for each $1\leq i<r$. If $I$ has linear quotients, then $I$ has linear resolution ([8, Proposition 8.2.1]).
@@
Let $G$ be a finite graph and fix a vertex $a$ of $G$. Let $G^{*}$ be a finite graph obtained by adding a new vertex $a^{*}$ to $G$ with
$E(G^{*})=E(G)\cup\{a^{*}b:ab\in E(G)\}.$
We say that $G^{*}$ is obtained from $G$ by duplicating a vertex $a$ of $G$. Note that $aa^{*}\not\in E(G^{*})$.
Proposition 2.2.
If $I(G)^{s}$ has linear quotients, then $I(G^{*})^{s}$ has linear quotients.

### v2

@@ Abstract.
Let $I(G)$ be the edge ideal of a gapfree graph $G$. An open conjecture of Nevo and Peeva states that $I(G)^{q}$ has linear resolution for $q\gg 0$. We present a promising approach to this challenging conjecture by investigating the stronger property of linear quotients. Specifically, we make the conjecture that if $I(G)^{q}$ has linear quotients for some integer $q\geq 1$, then $I(G)^{s}$ has linear quotients for all $s\geq q$. We give a partial solution to this conjecture, and identify conditions under which only finitely many powers need to be checked. It is known that if $G$ does not contain a cricket, a diamond, or a $C_{4}$, then $I(G)^{q}$ has linear resolution for $q\geq 2$. We construct a family of gapfree graphs $G$ containing cricket, diamond, $C_{4}$ together with $C_{5}$ as induced subgraphs of $G$ for which $I(G)^{q}$ has linear quotients for $q\geq 2$.
@@
edge ideal, gapfree graph, linear resolution, linear quotients, regularity.
@@ 1. Introduction
Let $G$ be a finite simple graph, i.e., a graph that does not contain loops or multiple edges, with vertex set $V(G)=\{1,\ldots,n\}$ and edge set $E(G)$. Let $S=K[x_{1},\ldots,x_{n}]$ denote the polynomial ring in $n$ variables over a field $K$. The edge ideal of $G$, denoted by $I(G)$, is the ideal of $S$ generated by squarefree quadratic monomials $x_{i}x_{j}$ for all $\{i,j\}\in E(G)$ (cf. [24]). The study of edge ideals has become a central topic in combinatorial commutative algebra, showcasing the deep interplay between combinatorial methods and algebraic structures; see, for instance, [3, 7, 8, 15, 16, 19, 20, 21, 22, 23] for various recent aspects of this study. The present paper builds upon this rich foundation, advancing the understanding of powers of edge ideals and their algebraic properties.
@@
Conjecture 1.1 has garnered significant interest (cf. [2, 6, 9, 10, 17] and references therein), but remains unresolved in general. While it has been verified for $q=2,3$ when $I(G)$ has small regularity (see [17]) and for random graphs with restricted edge probability (see [4]), progress has been limited. A promising approach to this challenging conjecture is to explore the stronger property of having linear quotients. This property has also been widely studied and, in particular, if a monomial ideal generated in a single degree has linear quotients then it has linear resolution (cf. [13]). Thus, as observed before, if $I(G)^{q}$ has linear quotients, for some $q\in{\mathbb{N}}$, then $G$ is necessarily a gapfree graph. We make the following conjecture:
@@
If $I(G)^{q}$ has linear quotients for some $q\in{\mathbb{N}}$, then $I(G)^{s}$ has linear quotients for all $s\geq q$.
@@
For instance, the pentagon $C_{5}$ is gapfree, but its complementary graph is not chordal. Consequently, $I(C_{5})$ does not have linear resolution. However, $C_{5}$ does not contain any induced subgraph that is isomorphic to a cricket (or a diamond, or $C_{4}$), so $I(C_{5})^{q}$ has linear resolution for all $q\geq 2$. On the other hand, it was shown (cf. [13, Corollary 10.1.7 and Theorems 10.1.9 and 10.2.5]) that if the complementary graph of $G$ is chordal then all powers of $I(G)$ have linear quotients. Beside the anticycles, as exhibited in a recent preprint [5] which appears after the first draft of this paper was written, there are currently no other known examples of a gapfree graph $G$ whose complementary graph is not chordal, yet for which all powers $I(G)^{q}$ with $q\geq 2$ have linear quotients. This motivates the search for natural classes of such graphs.
@@
In this paper, we establish a framework for addressing Conjecture 1.2. In Section 2, we examine the duplication technique, which when applied to graphs uses duplication of vertices in $G$ to produce new graphs $G^{x}$. This construction is standard in graph theory and has been used to obtain interesting algebraic results (cf. [11]). We show that for a general squarefree monomial ideal $I$ if a power $I^{q}$ has linear quotients, then so does $(I^{x})^{q}$ (Proposition 2.3). In Section 3, we show similar result holds for graphs when vertex duplication is replaced by graph expansion, where a vertex is replaced by a clique, as long as the resulting graph is still gapfree (Theorem 3.4).
@@
Section 4 studies Conjecture 4.1, which presents a strong case for Conjecture 1.2 by explicitly describing a possible linear quotient ordering for $I(G)^{s}$, for $s\geq q$, from that of $I(G)^{q}$. Particularly, Proposition 4.2 shows that Conjecture 4.1 holds for the pentagon $C_{5}$. In other words, all powers $I(C_{5})^{q}$ with $q\geq 2$ have linear quotients. Section 5 extends the study in Section 4 to a special class of graphs which are gapfree but may contain critical induced subgraphs such as crickets, diamonds and $C_{4}$. In Propositions 5.2 and 5.4, it is shown that, for these graphs, all powers $I(G)^{q}$ with $q\geq 2$ have linear quotients. The main result of the section (Theorem 5.7) extends the aforementioned result to an infinite family of graphs.
@@
Finally, Section 6 presents a refined analysis of modified versions of Conjectures 4.1 and 1.2, culminating in Theorem 6.4, which exhibits that if $I(G)^{2},\ldots,I(G)^{7}$ have linear quotients with a special ordering of their generators, then $I(G)^{q}$, for all $q\geq 2$, have linear quotients. This result reduces the verification of Conjectures 4.1 and 1.2 to finitely many powers and, therefore, suggests that resolving these conjectures may be within reach.
2. Duplicating a variable and linear quotients of powers
@@
Let $I$ be an ideal of $S$ which is generated by monomials of the same degree and let $\{u_{1},\ldots,u_{r}\}$ be the unique minimal monomial generators of $I$. We then call $u_{1},\ldots,u_{r}$ the generators of $I$ and use the notation $\mathcal{G}(I)$ to denote the set $\{u_{1},\ldots,u_{r}\}$. We say that $I$ has linear quotients if there is an ordering $u_{1}>\cdots>u_{r}$ of the generators of $I$ for which the colon ideal $(u_{1},\ldots,u_{i}):u_{i+1}$ is generated by a subset of $\{x_{1},\ldots,x_{n}\}$ for each $1\leq i<r$. If $I$ has linear quotients for $u_{1}>\cdots>u_{r}$, we call this ordering a linear quotients order for $I$. If $I$ has linear quotients, then $I$ has linear resolution ([13, Proposition 8.2.1]).
When showing an ideal has a linear quotients order, we will frequently examine individual monomial pairs. For simplicity, we will use the following notation below for two monomials $M$ and $N$
$M:N=\frac{M}{\gcd(M,N)}.$
In particular, $M\mid(M:N)N$ and $M:N$ is the minimal monomial under inclusion for which this property holds.
The following statement is an equivalent version of the definition of an ideal with linear quotients. The proof follows immediately from the definitions and so is not included.
Lemma 2.1 (Linear quotients).
Let $I$ be a monomial ideal, and fix an order $u_{1}>u_{2}>\cdots>u_{r}$ on the minimal generating set of $I$. Then the following statements are equivalent.
  1. (1)
    $I$ has linear quotients with respect to the order above;
  2. (2)
    if $\deg(u_{\alpha}:u_{\beta})>1$ for some indices $\alpha<\beta$, then there is some index $\gamma<\beta$ such that $u_{\gamma}:u_{\beta}=z$ where $z$ is a variable and $z\mid(u_{\alpha}:u_{\beta})$.
The equivalence of items $(1)$ and $(2)$ will be used throughout the paper when showing an ideal has linear quotients. The first result we show is that linear quotients are preserved under variable duplication. Since there are multiple ways to duplicate a variable, we establish a definition and notation.
Definition 2.2 (Duplicating a variable).
Let $I$ be a squarefree monomial ideal, suppose $x$ is a variable, and suppose $m_{1},\ldots,m_{q}$ are all the monomials in ${\mathcal{G}}(I)$ that are divisible by $x$. Let $y$ be a new variable that does not divide any generator of $I$. The duplicate ideal of $I$ by $x$ is the ideal
$I^{x}=I+\big{(}\frac{m_{1}}{x}\cdot y,\ldots,\frac{m_{q}}{x}\cdot y\big{)}.$
If $x$ does not divide any generator of $I$, then $I^{x}=I$.
We next show that linear quotients of squarefree monomial ideals are preserved under variable duplication. For a monomial $M$ and a variable $x$, we use the notation $\deg_{x}(M)$ to denote the maximum value of $i$ such that $x^{i}\mid M$.
Proposition 2.3.
Let $I$ be an equigenerated squarefree monomial ideal, $x$ a variable, and $s$ a positivie integer. If $I^{s}$ has linear quotients, then $(I^{x})^{s}$ has linear quotients.
Proof.
Fix a variable $x$. If $x$ does not appear in ${\mathcal{G}}(I)$ then $I^{x}=I$ and we are done by assumption, so we assume $x$ divides a generator of $I$. Let $y$ be the duplicate of $x$ in $I^{x}$. Note that, for each $u\in{\mathcal{G}}(I^{s})$ with $\deg_{x}(u)=d>0$, the following $d$ monomials obtained from $u$
(1) $\frac{\,u\,}{x}\cdot y,\frac{\,u\,}{x^{2}}\cdot y^{2},\dots,\frac{\,u\,}{x^{d}}\cdot y^{d}$
are the elements of ${\mathcal{G}}((I^{x})^{s})$ which are not in ${\mathcal{G}}(I^{s})$. Order the generators of $(I^{x})^{s}$ by extending a linear quotients order of $I^{s}$ in the following way: If $u\in{\mathcal{G}}(I^{s})$ with $\deg_{x}(u)>0$, insert the new generators obtained from $u$ immediately following $u$ in the order they appear in (1).
Let $w_{1}>\cdots>w_{q}$ denote the generators of $(I^{x})^{s}$ in this order. We claim that this is a linear quotients order of $(I^{x})^{s}$. Let $t\geq 2$ and
$J:=(w_{1},\dots,w_{t-1}):w_{t}.$
Consider $w_{\ell}:w_{t}$ for $\ell<t$.
CASE 1:
Suppose $w_{t}\in{\mathcal{G}}(I^{s})$, i.e., $w_{t}$ is not divisible by $y$.
$\cdot$ If $w_{\ell}\in{\mathcal{G}}(I^{s})$, we are done since $I^{s}$ has linear quotients.
$\cdot$ If $w_{\ell}\notin{\mathcal{G}}(I^{s})$, then $w_{\ell}$ is divisible by $y$, which implies that $w_{\ell}:w_{t}\in(y)$.
    If $x\in{\mathcal{G}}(J)$, there exists $w_{j}\in{\mathcal{G}}(I^{s})$ with $j<t$ such that $w_{j}:w_{t}=x$. Let $w_{k}=w_{j}y/x$. Notice that $j<k<t$ by the definition of the order on ${\mathcal{G}}((I^{x})^{s})$. Thus, we have $w_{k}:w_{t}=y\in{\mathcal{G}}(J)$, as desired.
    If $x\notin{\mathcal{G}}(J)$, then let $\deg_{x}(w_{\ell})=\alpha_{x}\geq 0$ and $\deg_{y}(w_{\ell})=\alpha_{y}\geq 1$. Then $w_{\ell}=x^{\alpha_{x}}y^{\alpha_{y}}M$ for a monomial $M$. Consider the monomial $w_{k}=x^{\alpha_{x}+\alpha_{y}}M\in{\mathcal{G}}(I^{s})$ and notice that $k<{\ell}$ by the definition of the order. Since $I^{s}$ has linear quotients, there exists $w_{j}\in{\mathcal{G}}(I^{s})$ with $j<t$ and a variable $z$ such that $z=(w_{j}:w_{t})$ and $z\mid(w_{k}:w_{t})$. Note that $z\neq x$ since $x\notin{\mathcal{G}}(J)$, and $z\neq y$ since $w_{j},w_{t}\in{\mathcal{G}}(I^{s})$. Therefore, $w_{\ell}:w_{t}$ is divisible by $z$, and hence $J$ is linear.
CASE 2:
Suppose $w_{t}\notin{\mathcal{G}}(I^{s})$, i.e., $w_{t}$ is divisible by $y$.
Our first goal is to show $x\in{\mathcal{G}}(J)$. Set $w_{j}=w_{t}x/y$. Then $j<t$ by the definition of the order. Thus, $w_{j}:w_{t}=x\in{\mathcal{G}}(J)$.
Let $w_{\ell}=x^{\alpha_{x}}y^{\alpha_{y}}M$ for a monomial $M$ where $\alpha_{x}=\deg_{x}(w_{\ell})$ and $\alpha_{y}=\deg_{y}(w_{\ell})$. Similarly, let $w_{t}=x^{\beta_{x}}y^{\beta_{y}}N$ for a monomial $N$ where $\beta_{x}=\deg_{x}(w_{t})$ and $\beta_{y}=\deg_{y}(w_{t})\geq 1$. Consider the monomials $w_{L}=x^{\alpha_{x}+\alpha_{y}}M\text{ and }w_{T}=x^{\beta_{x}+\beta_{y}}N$ in ${\mathcal{G}}(I^{s})$ where $L\leq{\ell}$ and $T<t$ by the definition of the order.
If $\alpha_{x}>\beta_{x}$, then $x\in{\mathcal{G}}(J)$ divides $w_{\ell}:w_{t}$, and we are done. Suppose $\alpha_{x}\leq\beta_{x}$. This implies that ${\ell}<T$ by the definition of the order. Since $I^{s}$ has linear quotients and $L<T$, there exists $w_{j}\in{\mathcal{G}}(I^{s})$ with $j<T$ such that $w_{j}:w_{T}=z$ for some variable $z$ dividing $w_{L}:w_{T}$. Note that $z\neq y$. Let $\alpha=\alpha_{x}+\alpha_{y}$ and $\beta=\beta_{x}+\beta_{y}$.
Case 2.1: Suppose $\alpha\leq\beta$. In this case, $w_{L}:w_{T}=M:N$ which implies that $w_{\ell}:w_{t}$ is divisible by $z$. It remains to show $z\in{\mathcal{G}}(J)$. Since $w_{j}:w_{T}=z$, the monomial $w_{j}$ has one of the following forms:
  1. (a)
    $w_{j}=x^{\beta-1}zN$ or
  2. (b)
    $w_{j}=x^{\beta}Nz/z^{\prime}$ for a variable $z^{\prime}$ dividing $N$ where $z^{\prime}\neq z$.
If (a) holds, let $w_{k}=x^{\beta_{x}}y^{\beta_{y}-1}zN$ where $j\leq k<T<t$ by the definition of the order. Thus, $w_{k}:w_{t}=z\in{\mathcal{G}}(J)$, as desired. Similarly, if (b) holds, let $w_{k}=x^{\beta_{x}}y^{\beta_{y}}Nz/z^{\prime}$ where $j<k<T<t$ by the definition of the order. Hence, $w_{k}:w_{t}=z\in{\mathcal{G}}(J)$.
Case 2.2: Suppose $\alpha>\beta$. It then follows from the assumption $\alpha_{x}\leq\beta_{x}$ that $\alpha_{y}>\beta_{y}$. In this case, $w_{\ell}:w_{t}=y^{\alpha_{y}-\beta_{y}}(M:N)$ and $w_{L}:w_{T}=x^{\alpha-\beta}(M:N)$. So, $w_{\ell}:w_{t}$ is divisible by both $y$ and $z$. It remains to show either $y\in{\mathcal{G}}(J)$ or $z\in{\mathcal{G}}(J)$. Notice that it is possible to have $z=x$.
Suppose $z=x$. Then $w_{j}=x^{\beta+1}N/z^{\prime}$ for some variable $z^{\prime}$ dividing $N$ since $w_{j}:w_{T}=z$. Let $w_{k}=x^{\beta_{x}}y^{\beta_{y}+1}N/z^{\prime}$. Note that $j<k<T<t$ by the definition of the order and $w_{k}:w_{t}=y\in{\mathcal{G}}(J)$.
Suppose $z\neq x$. This implies that $z$ divides $M:N$. In this case, $w_{j}$ is in one of the forms described in (a) or (b) from Case 2.1. The remainder of the proof follows similarly and it results with $z\in{\mathcal{G}}(J)$. ∎
3. Expansion, and linear quotients of powers of edge ideals
For the remainder of the paper, we focus our attention on edge ideals of graphs. Let $G$ be a finite graph on $V(G)=\{1,\ldots,n\}$ with $E(G)=\{e_{1},\ldots,e_{s}\}$, and let $S=K[x_{1},\ldots,x_{n}]$ be the polynomial ring in $n$ variables over a field $K$. For an edge $e=\{i,j\}\in E(G)$, we define $u_{e}=x_{i}x_{j}\in S$. Recall that the edge ideal of $G$ is the ideal $I(G)$ of $S$ which is generated by those monomials $u_{e}$ with $e\in E(G)$. To simplify the notation, unless misleading, we write $e$ instead of $u_{e}$. For example, $e_{1}^{3},e_{1}e_{2}^{2}\in I(G)^{3}$. In addition, we may write $ij\in E(G)$ instead of $\{i,j\}\in E(G)$. Lastly, we may refer to and use the variables $x_{1},\ldots,x_{n}$ as vertices of $G$.
@@ 3. Expansion, and linear quotients of powers of edge ideals
Let $G$ be a finite graph and fix a vertex $x$ of $G$.
  • The graph $G^{x}$, called the duplication of $G$ at $x$, is a finite graph obtained by adding a new vertex $y$ to $G$ with
    $E(G^{x})=E(G)\cup\{yb:xb\in E(G)\}.$
    We say that $G^{x}$ is obtained from $G$ by duplicating the vertex $x$ of $G$. Note that $xy\not\in E(G^{x})$, and also $I(G^{x})=I(G)^{x}$ in the sense of 2.2. We call the new vertex $y$ the duplicate of $x$ in $G^{x}$.
  • The graph $G^{[x]}$, called the expansion of $G$ at $x$, is the graph with vertex set $V(G^{[x]})=V(G^{x})$ and $E(G^{[x]})=E(G^{x})\cup\{xy\}$.
  • For any $W\subseteq V(G^{[x]})=V$ and any monomial $m\in I(G^{[x]})^{s}$, define $m_{W}$ to be the monomial formed by localizing $m$ at $W$. That is, if $m=\prod_{v_{i}\in V}v_{i}^{s_{i}}$, then $m_{W}=\prod_{v_{i}\in W}v_{i}^{s_{i}}$.
The following example illustrates the definitions above.
Example 3.2.
Let $G$ be the graph $C_{4}$ labeled as in the left hand graph below. Then $G^{x}$ is the central graph and $G^{[x]}$ is the graph on the right. Consider the monomial $m=(xy)^{2}(xa)(bc)\in I(G^{[s]})^{4}$. If $W=\{x,c\}$, then $m_{W}=x^{3}c.$
$x$$a$$c$$b$ $x$$y$$a$$c$$b$ $x$$y$$a$$c$$b$
Figure 2. Graph $G$, $G^{x}$ and $G^{[x]}$
We state some useful facts regarding the duplicated vertex $x$ and its duplication $y$. The proofs are straightforward from the definitions and so are not included. First, recall that
$N_{G}(x)=\{z\in V(G):xz\in E(G)\}\quad\mbox{and}\quad N_{G}[x]=N_{G}(x)\cup\{x\}.$
Lemma 3.3.
Let $G$ be a gapfree graph. Using the notation of 3.1,
  • $N_{G^{x}}(x)=N_{G^{x}}(y)=N_{G}(x)=N_{G^{[x]}}\setminus\{y\}$.
  • $N_{G^{[x]}}[x]=N_{G^{[x]}}[y]$.
  • $G^{[x]}$ is gapfree if and only if $V(G)\setminus N_{G}[x]$ is an independent set.
We are now ready to prove the main theorem of this section.
Theorem 3.4.
Let $G$ be a graph such that $I(G)^{s}$ has linear quotients for some $s\geq 1$. Fix $x\in V(G)$. Then $I(G^{[x]})^{s}$ has linear quotients if and only if $G^{[x]}$ is gapfree.